# Optimal and fast online change point estimation in linear regression

### v1

@@ 5 Proofs
using the inequality $1/\log^{2/3}n\leq 1$ for $n>2$. Thus, the statement for a kink is true with $r^{*}_{K}=9\left(300/\delta_{0}^{2}\right)^{2/3}+1$. ∎
@@
We fix $\alpha_{-}=\beta_{-}=0$ and some $\alpha_{+}$ and $\beta_{+}$, because a lower rate obtained over the fixed parameters is also a lower rate for the whole parameter space. We set $M\coloneqq\left\lfloor\left(1-2\delta_{0}\right)/3b\psi_{n}\right\rfloor$, with $b$ a positive constant independent of $n$, to be determined later. We choose $t_{0},...,t_{M}\in(\delta_{0},1-\delta_{0})$ such that $t_{j}-t_{j-1}=3b\psi_{n}$ for $j=0,...,M$ and $M>1$. As $\frac{1}{n}\leq\psi_{n}\ll 1$, this is possible. Now we assume that
@@
Let $\mathcal{B}_{j}\coloneqq\left\{\left|\tilde{\tau}_{n}-t_{j}\right|\leq b\psi_{n}\right\}$. The distance between two distinct points of $t_{0},...,t_{M}$ is always greater than $3b\psi_{n}$, so the events $\mathcal{B}_{j}$ are all mutually exclusive. Note also that $\bigcup_{j=0}^{M-1}\mathcal{B}_{j}\subset\left\{\left|\tilde{\tau}-t_{M}\right|>b\psi_{n}\right\}$. We denote
@@
the likelihood function of $X_{1},...,X_{n}$ under the change point $t_{j}$. Next we can calculate
@@
$\qquad\leq\mathbb{P}_{t_{M}}\left(\left|\tilde{\tau}_{n}-t_{M}\right|>b\psi_{n}\right)\leq\max_{0\leq j\leq M}\mathbb{P}_{t_{j}}\left(\left|\tilde{\tau}_{n}-t_{j}\right|>b\psi_{n}\right).$
@@
$\frac{\mathbf{d}\mathbb{P}_{t_{M}}}{\mathbf{d}\mathbb{P}_{t_{j}}}$ $=\frac{\prod_{i=1}^{n}\exp\left\{-\frac{1}{2}\left(X_{i}-f_{M}(i)\right)^{2}\right\}}{\prod_{i=1}^{n}\exp\left\{-\frac{1}{2}\left(X_{i}-f_{j}(i)\right)^{2}\right\}}$
$=\exp\left\{\sum_{i=\left\lfloor t_{j}n\right\rfloor+1}^{n}\frac{1}{2}\left(\left(X_{i}-f_{j}(i)\right)^{2}-\left(X_{i}-f_{M}(i)\right)^{2}\right)\right\},$
and
$\frac{1}{2}\left(\left(X_{i}-f_{j}(i)\right)^{2}-\left(X_{i}-f_{M}(i)\right)^{2}\right)$ $=\frac{1}{2}\left(2X_{i}\left(f_{M}(i)-f_{j}(i)\right)+f_{j}^{2}(i)-f_{M}^{2}(i)\right)$
@@
where $\varepsilon_{i}=X_{i}-f_{j}(i)\sim\mathcal{N}(0,1)$ if $\left(X_{i}\right)_{i=1}^{n}\sim\mathbb{P}_{t_{j}}$. Now we can use the moment generating function of the standard normal, $\mathbb{E}\left[\exp\left\{tZ\right\}\right]=\exp\left\{\frac{1}{2}t^{2}\right\}$, to get
@@
Note that $g_{n}$ is not dependent on $j$.
Next we define $Z_{j}$ as
@@
Since $\tilde{\tau}$ is a Markov stopping time, the event $\mathcal{B}_{j}=\left\{\left|\tilde{\tau}-t_{j}\right|\leq b\psi_{n}\right\}$ is $\mathcal{F}_{u_{j}}$-measurable and is therefore independent of $X_{i}$ and $\varepsilon_{i}$, for $i=\left\lfloor u_{j}n\right\rfloor+1,...,n$. We can use this independence and (19), to rewrite the sum in (18) as
@@
$\qquad=\mathbb{E}_{t_{j}}\left[\exp\left\{\sum_{i=\left\lfloor t_{j}n\right\rfloor+1}^{\left\lfloor u_{j}n\right\rfloor}\varepsilon_{i}\left(f_{M}(i)-f_{j}(i)\right)-\sum_{i=\left\lfloor t_{j}n\right\rfloor+1}^{\left\lfloor u_{j}n\right\rfloor}\frac{1}{2}\left(f_{M}(i)-f_{j}(i)\right)^{2}\right\}\mathbb{I}\left(\mathcal{B}_{j}\right)\right]$
@@
$\qquad=\mathbb{E}_{t_{j}}\left[\exp\left\{\left(-\sum_{i=\left\lfloor t_{j}n\right\rfloor+1}^{\left\lfloor u_{j}n\right\rfloor}\varepsilon_{i}f_{j}(i)\right)-g_{n}\right\}\mathbb{I}\left(\mathcal{B}_{j}\right)\right]$
$\qquad=\mathbb{E}_{t_{j}}\left[\exp\left\{\sqrt{\sum_{i=\left\lfloor t_{j}n\right\rfloor+1}^{\left\lfloor u_{j}n\right\rfloor}\left(f_{j}(i)\right)^{2}}Z_{j}-g_{n}\right\}\mathbb{I}\left(\mathcal{B}_{j}\right)\right]$
$\qquad\geq\mathbb{E}_{t_{j}}\left[\exp\left\{\sqrt{\sum_{i=\left\lfloor t_{j}n+1\right\rfloor}^{\left\lfloor u_{j}n\right\rfloor}\left(f_{j}(i)\right)^{2}}Z_{j}-g_{n}\right\}\mathbb{I}\left(\mathcal{B}_{j}\right)\mathbb{I}(Z_{j}\geq 0)\right]$
$\qquad\geq\exp\left\{-g_{n}\right\}\mathbb{P}_{t_{j}}\left(\mathcal{B}_{j}\cap\left\{Z_{j}\geq 0\right\}\right).$
@@
$\mathbb{P}\left(\mathcal{B}_{j}\cap\left\{Z_{j}\geq 0\right\}\right)$ $=\mathbb{P}(\mathcal{B}_{j})+\mathbb{P}(Z_{j}\geq 0)-\mathbb{P}_{t_{j}}(\mathcal{B}_{j}\cup\left\{Z_{j}\geq 0\right\})$
$\geq\frac{3}{4}+\frac{1}{2}-1=\frac{1}{4}.$
@@
$\frac{n}{\log n}\exp\left\{-\frac{\alpha_{+}^{2}b}{2}\log n\right\}$ $=\frac{n}{\log n}\exp\left\{-\frac{1}{2}\log n\right\}=\frac{\sqrt{n}}{\log n}\to\infty\text{, for }n\to\infty.$
Thus, $\frac{n}{\log n}\exp\left\{-g_{n}\right\}\to\infty$ for $n\to\infty$, which is a contradiction to (20), so the statement is true.
@@
Thus, $\left({n^{1/3}}\middle/{\log^{{1}/{3}}}\right)\exp\left\{-g_{n}\right\}\to\infty$, for $n\to\infty$, which is a contradiction to (20). ∎

### v2

@@ 5 Proofs
by the fact $1/\log^{2/3}n\leq 1$ for $n>2$. Thus, the statement for a kink holds with $r^{*}_{K}=9\left(300/\delta_{0}^{2}\right)^{2/3}+1$.
Analogous to the jump case, the requirement on the size of historical data can be relaxed from $k\asymp n$ to $k\asymp n^{8/9}\log(n)^{1/9}$; thanks to (15), this relaxed choice still guarantees (23). ∎
Proof of Proposition 2.
Note that $N_{J}\coloneqq b_{J}\log n$, $M_{J}\coloneqq 2N_{J}+(m\mod N_{J})$, $b_{J}\coloneqq 10^{3}/\left(2c^{2}\right)$ and $\rho_{J}\coloneqq 4c/5$ in the jump case, while $N_{K}\coloneqq b_{k}n^{2/3}\log^{1/3}n$, $M_{K}\coloneqq 2N_{K}+(m\mod N_{K})$, $b_{K}\coloneqq(300/c^{2})^{1/3}$ and $\rho_{K}=4c/(5n)$ in the kink case. We set $Z_{J,m}\coloneqq\sum_{i=1}^{M}\varepsilon_{m-M+i}/\sqrt{M}\sim\mathcal{N}(0,1)$ and $Z_{k,m}\coloneqq(1/\sqrt{d_{M}})\sum_{i=1}^{M}i\varepsilon_{m-M+i}$. We define events $\mathcal{A}_{J}=\mathcal{A}$ in (9) and $\mathcal{A}_{K}=\mathcal{A}$ in (19). In the proof of Theorem 1 we showed that $\mathbb{P}\left(\mathcal{A}^{c}_{J}\right)\leq n^{-4}$ and $\mathbb{P}\left(\mathcal{A}^{c}_{K}\right)\leq n^{-3}$, and thus $\mathbb{P}\left(\left(\mathcal{A}_{J}\cap\mathcal{A}_{K}\right)^{c}\right)\ll n^{-3}$. Also the corresponding detector will not make a false detection under the events $\mathcal{A}_{J}$ and $\mathcal{A}_{K}$.
For a jump
Fix $\theta_{0}\in\Theta_{\delta_{0}}^{J}$. We only need to show $\{\hat{\tau}_{K,n}<\hat{\tau}_{J,n}\}\subseteq\{\mathcal{A}_{J}\cap\mathcal{A}_{K}\}^{c}$. Note that we already showed in the proof of Theorem 1 that under a jump and the event $\mathcal{A}_{J}$ is $\hat{\tau}_{J,n}\leq\tau_{0}+3N_{J}/n$ and that under the event $\mathcal{A}_{K}$ is $\tau_{0}\leq\hat{\tau}_{K,n}$ Now we choose $m\in\left[\tau_{0}n,\tau_{0}n+3N_{J}\right]$:
$K_{m}$ $=\left|\frac{1}{d_{M_{K}}}\sum_{i=1}^{M_{K}}i\varepsilon_{m-M_{K}+i}+\frac{1}{d_{M_{K}}}\sum_{i=1}^{M_{K}}i\left(f_{+}\left(\frac{m-M_{K}+i}{n}\right)-\hat{f}_{-}\left(\frac{m-M_{K}+i}{n}\right)\right)\right|$
$\leq\left|\frac{1}{d_{M_{K}}}\sum_{i=1}^{M_{K}}i\left(f_{+}\left(\frac{m-M_{K}+i}{n}\right)-f_{-}\left(\frac{m-M_{K}+i}{n}\right)\right)\mathbb{I}\left(i>\tau_{0}n-m+M_{K}\right)\right|$ (24a)
$\qquad+\left|\frac{1}{d_{M_{K}}}\sum_{i=1}^{M_{K}}i\varepsilon_{m-M_{K}+i}\right|$ (24b)
$\qquad\qquad+\left|\frac{1}{d_{M_{K}}}\sum_{i=1}^{M_{K}}i\left(f_{-}\left(\frac{m-M_{K}+i}{n}\right)-\hat{f}_{-}\left(\frac{(m-M_{K}+i}{n}\right)\right)\right|,$ (24c)
where both terms in (24b) and (24c) can be bounded by $c/(10n)$, see (20) and (21). For (24a) we get
$\left|\frac{1}{d_{M_{K}}}\sum_{i=1}^{M_{K}}i\left(f_{+}\left(\frac{m-M_{K}+i}{n}\right)-f_{-}\left(\frac{m-M_{K}+i}{n}\right)\right)\mathbb{I}\left(i>\tau_{0}n-m+M_{K}\right)\right|$
$=\,$ $\left|\frac{1}{d_{M_{k}}}\sum_{i=\tau_{0}n-m+M_{K}}^{M_{K}}i\left(\left(\beta_{+}-\beta_{-}\right)\left(\frac{m-M_{K}+i}{n}-\tau_{0}\right)+\alpha_{+}-\alpha_{-}\right)\right|$
$\leq\,$ $\frac{1}{d_{M_{K}}}\sum_{i=0}^{m-\tau_{0}n}\left(i^{2}\frac{\left|\beta_{+}-\beta_{-}\right|}{n}+i\left|\alpha_{+}-\alpha_{-}\right|\right)\leq\frac{1}{d_{M_{K}}}\sum_{i=0}^{3N_{J}}\left(i^{2}\frac{\left|\beta_{+}-\beta_{-}\right|}{n}+i\left|\alpha_{+}-\alpha_{-}\right|\right)$
$\leq\,$ $\frac{3N_{J}(3N_{J}+1)(6N_{J}+1)}{3N_{K}(3N_{K}+1)(6N_{K}+1)}\frac{\left|\beta_{+}-\beta_{-}\right|}{n}+\frac{9N_{J}(N_{J}+1)}{3N_{K}(3N_{K}+1)(6N_{K}+1)}\left|\alpha_{+}-\alpha_{-}\right|$
$\asymp\,$ $\frac{\log^{2}n}{n^{3}}\left|\beta_{+}-\beta_{-}\right|+\frac{\log n}{n^{2}}\left|\alpha_{+}-\alpha_{-}\right|.$
Now we can bound (24a) by c/(5n){c}/{(5n)} for $n$ large enough and finally get that $K_{m}\leq c/(5n)+c/(10n)+c/(10n)=2/(5n)<\rho_{K}$. Hence we get that under both events $\mathcal{A}_{J}$ and $\mathcal{A}_{K}$ is $\hat{\tau}_{k,n}>\tau_{0}+3N_{J}/n\geq\hat{\tau}_{J,n}$.
For a kink
For $\theta_{0}\in\Theta^{K}_{\delta_{0}}$, we want to show $\{\hat{\tau}_{J,n}<\hat{\tau}_{K,n}\}\subseteq\{\mathcal{A}_{J}\cap\mathcal{A}_{K}\}^{c}$. In the proof of Theorem 1 we showed that $\hat{\tau}_{K,n}\leq\tau_{0}+3N_{K}/n$ in this case under the event $\mathcal{A}_{K}$. For $m\in\left[\tau_{0}n,\tau_{0}n+3N_{K}\right]$,
$J_{m}$ $=\left|\frac{1}{M_{J}}\sum_{i=1}^{M_{J}}\varepsilon_{m-M_{J}+i}+\frac{1}{M_{J}}\sum_{i=1}^{M_{J}}\left(f_{+}\left(\frac{m-M_{J}+i}{n}\right)-\hat{f}_{-}\left(\frac{m-M_{J}+i}{n}\right)\right)\right|$
$\leq\left|\frac{1}{M_{J}}\sum_{i=1}^{M_{J}}\left(f_{+}\left(\frac{m-M_{J}+i}{n}\right)-f_{-}\left(\frac{m-M_{J}+i}{n}\right)\right)\mathbb{I}\left(i>\tau_{0}n-m+M_{J}\right)\right|$ (25a)
$\qquad+\left|\frac{1}{M_{J}}\sum_{i=1}^{M_{J}}\varepsilon_{m-M_{J}+i}\right|$ (25b)
$\qquad\qquad+\left|\frac{1}{M_{J}}\sum_{i=1}^{M_{J}}\left(f_{-}\left(\frac{m-M_{J}+i}{n}\right)-\hat{f}_{-}\left(\frac{m-M_{J}+i}{n}\right)\right)\right|.$ (25c)
We have shown in (10) that under the event $\mathcal{A}_{J}$ (25b) is bounded by $c/10$ and in (11) that (25c) is bounded by $c/10$. For the part (25a) we get
$\left|\frac{1}{M_{J}}\sum_{i=1}^{M_{J}}\left(f_{+}\left(\frac{m-M_{J}+i}{n}\right)-f_{-}\left(\frac{m-M_{J}+i}{n}\right)\right)\mathbb{I}\left(i>\tau_{0}n-m+M_{J}\right)\right|$
$\leq\,$ $\frac{1}{M_{J}}\sum_{i=1}^{M_{J}}\left|\beta_{+}-\beta_{-}\right|\left(\frac{m-M_{J}+i}{n}-\tau_{0}\right)+\left|\alpha_{+}-\alpha_{-}\right|$
$\leq\,$ $\frac{1}{M_{J}}\sum_{i=1}^{M_{J}}\left|\beta_{+}-\beta_{-}\right|\left(\frac{3N_{K}-M_{J}+i}{n}\right)+\left|\alpha_{+}-\alpha_{-}\right|\leq\left|\beta_{+}-\beta_{-}\right|\frac{3N_{K}}{n}+\left|\alpha_{+}-\alpha_{-}\right|$
$=\,$ $\left|\beta_{+}-\beta_{-}\right|\frac{3b_{k}\log^{1/3}n}{n^{1/3}}+\left|\alpha_{+}-\alpha_{-}\right|<\frac{c}{5},$
for $n$ large enough, since we have that $\left|\alpha_{+}-\alpha_{-}\right|<\delta_{0}$. As a result we get $J_{m}<c/10+c/10+c/5=2c/5<\rho_{J}$ and now have under the event $\left\{\mathcal{A}_{J}\cap\mathcal{A}_{K}\right\}$ that $\hat{\tau}_{J,n}>\tau_{0}+3N_{K}/n\geq\hat{\tau}_{K,n}$. ∎
Proof of Corollary 3.
We prove the bound on the expected detection delay by utilizing (2) and (3). By these equations we get for any detector $\tau_{n}$ the inequality
$\bigl{(}\mathbb{E}_{\tau}\left[\hat{\tau}_{n}-\tau\mid\hat{\tau}_{n}\geq\tau\right]\bigr{)}^{2}\leq\frac{\mathbb{E}_{\tau}\left[\left(\hat{\tau}_{n}-\tau\right)^{2}\right]}{\bigl{(}1-\mathbb{P}_{\tau}(\hat{\tau}_{n}<\tau)\bigr{)}}.$ (26)
In the proof of Theorem 1 we showed that $\mathbb{P}_{\tau}\left(\hat{\tau}_{J,n}<\tau\right)\leq n^{-3}$ and
$\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{J}_{\delta_{0}}}\mathbb{E}_{\tau}\left[\left(\frac{n}{\log n}\left(\hat{\tau}_{J,n}-\tau\right)\right)^{2}\right]\;\leq\;r^{*}_{J}<\infty.$
By plugging this into (26) we get for $n$ large enough a bound on the expected detection delay
$\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{J}_{\delta_{0}}}\frac{n}{\log n}\mathbb{E}_{\tau}\left[\hat{\tau}_{J,n}-\tau\mid\hat{\tau}_{J_{n}}\geq\tau\right]\leq\sqrt{r_{J}^{*}}.$
To bound the average run length, we recall that $\hat{\tau}_{J,n}$ is set to $1$ if no detection is made. Thus the bound on the average run length under the null is
$\mathbb{E}_{0}\left[\hat{\tau}_{J,n}\right]\geq\mathbb{P}_{0}(\hat{\tau}_{J,n}=1)\geq 1-n^{-3}.$
The proof for the kink case is analogous. We have already shown that $\mathbb{P}\left(\hat{\tau}_{K,n}\leq\tau\right)\leq n^{-4}$ and
$\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{K}_{\delta_{0}}}\mathbb{E}_{\tau}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\left(\hat{\tau}_{K,n}-\tau\right)\right)^{2}\right]\leq r^{*}_{K}<\infty.$
With these results we bound the expected detection delay by
$\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{K}_{\delta_{0}}}\frac{n^{1/3}}{\log^{1/3}n}\mathbb{E}_{\tau}\left[\hat{\tau}_{K,n}-\tau|\hat{\tau}_{K,n}\geq\tau\right]\leq\sqrt{r_{K}^{*}},$
and the average run length by
$\mathbb{E}_{0}\left[\hat{\tau}_{K,n}\right]\geq 1-n^{-4}.$
Combining the above statements shows the claim for the FLOC detector $\hat{\tau}_{n}=\min\{\hat{\tau}_{J,n},\hat{\tau}_{K,n}\}$. ∎
@@
We fix $\alpha_{-}=\beta_{-}=0$ and some $\alpha_{+}$ and $\beta_{+}$, this is possible, since a lower rate obtained over the fixed parameters is also a lower rate for the whole parameter space. We set $M\coloneqq\left\lfloor\left(1-2\delta_{0}\right)/(3b\psi_{n})\right\rfloor$, with $b$ a positive constant independent of $n$, to be determined later. We choose $t_{0},\ldots,t_{M}\in(\delta_{0},1-\delta_{0})$ such that $t_{j}-t_{j-1}=3b\psi_{n}$ for $j=0,\ldots,M$ and $M>1$. As $\frac{1}{n}\leq\psi_{n}\ll 1$, this is possible. Now we assume that
@@
Let $\mathcal{B}_{j}\coloneqq\left\{\left|\tilde{\tau}_{n}-t_{j}\right|\leq b\psi_{n}\right\}$. The distance between any two points of $t_{0},\ldots,t_{M}$ is greater than $3b\psi_{n}$, so the events $\mathcal{B}_{j}$ are all mutually exclusive. Note also that $\bigcup_{j=0}^{M-1}\mathcal{B}_{j}\subset\left\{\left|\tilde{\tau}-t_{M}\right|>b\psi_{n}\right\}$. We denote
@@
the likelihood function of $X_{1},\ldots,X_{n}$ under the change point $t_{j}$. Next we can calculate
@@
$\leq\,$ $\mathbb{P}_{t_{M}}\left(\left|\tilde{\tau}_{n}-t_{M}\right|>b\psi_{n}\right)\leq\max_{0\leq j\leq M}\mathbb{P}_{t_{j}}\left(\left|\tilde{\tau}_{n}-t_{j}\right|>b\psi_{n}\right).$
@@
$\frac{\mathbf{d}\mathbb{P}_{t_{M}}}{\mathbf{d}\mathbb{P}_{t_{j}}}=\frac{\prod_{i=1}^{n}\exp\left\{-\frac{1}{2}\left(X_{i}-f_{M}(i)\right)^{2}\right\}}{\prod_{i=1}^{n}\exp\left\{-\frac{1}{2}\left(X_{i}-f_{j}(i)\right)^{2}\right\}}$ $=\exp\left\{\sum_{i=\left\lfloor t_{j}n\right\rfloor+1}^{n}\frac{1}{2}\left(\left(X_{i}-f_{j}(i)\right)^{2}-\left(X_{i}-f_{M}(i)\right)^{2}\right)\right\},$
$\text{and}\qquad\frac{1}{2}\left(\left(X_{i}-f_{j}(i)\right)^{2}-\left(X_{i}-f_{M}(i)\right)^{2}\right)$ $=\frac{1}{2}\left(2X_{i}\left(f_{M}(i)-f_{j}(i)\right)+f_{j}^{2}(i)-f_{M}^{2}(i)\right)$
@@
where $\varepsilon_{i}=X_{i}-f_{j}(i)\sim\mathcal{N}(0,1)$ if $\left(X_{i}\right)_{i=1}^{n}\sim\mathbb{P}_{t_{j}}$. Now we can use the moment generating function of the standard normal distribution, $\mathbb{E}\left[\exp\left\{tZ\right\}\right]=\exp\left\{t^{2}/2\right\}$, to get
@@
Note that $g_{n}$ is not dependent on $j$. Next we define $Z_{j}$ as
@@
As $\tilde{\tau}$ is a Markov stopping time, $\mathcal{B}_{j}=\left\{\left|\tilde{\tau}-t_{j}\right|\leq b\psi_{n}\right\}$ is $\mathcal{F}_{u_{j}}$-measurable and is thus independent of $X_{i}$ and $\varepsilon_{i}$, for $i\in\{\left\lfloor u_{j}n\right\rfloor+1,\ldots,n\}$. By this independence and (30), we rewrite the sum in (29) as
@@
$=\mathbb{E}_{t_{j}}\left[\exp\left\{\sum_{i=\left\lfloor t_{j}n\right\rfloor+1}^{\left\lfloor u_{j}n\right\rfloor}\varepsilon_{i}\left(f_{M}(i)-f_{j}(i)\right)-\sum_{i=\left\lfloor t_{j}n\right\rfloor+1}^{\left\lfloor u_{j}n\right\rfloor}\frac{1}{2}\left(f_{M}(i)-f_{j}(i)\right)^{2}\right\}\mathbb{I}\left(\mathcal{B}_{j}\right)\right]$
@@
$=\mathbb{E}_{t_{j}}\left[\exp\left\{\left(-\sum_{i=\left\lfloor t_{j}n\right\rfloor+1}^{\left\lfloor u_{j}n\right\rfloor}\varepsilon_{i}f_{j}(i)\right)-g_{n}\right\}\mathbb{I}\left(\mathcal{B}_{j}\right)\right]$
$=\mathbb{E}_{t_{j}}\left[\exp\left\{\sqrt{\sum_{i=\left\lfloor t_{j}n\right\rfloor+1}^{\left\lfloor u_{j}n\right\rfloor}\left(f_{j}(i)\right)^{2}}Z_{j}-g_{n}\right\}\mathbb{I}\left(\mathcal{B}_{j}\right)\right]$
$\geq\mathbb{E}_{t_{j}}\left[\exp\left\{\sqrt{\sum_{i=\left\lfloor t_{j}n+1\right\rfloor}^{\left\lfloor u_{j}n\right\rfloor}\left(f_{j}(i)\right)^{2}}Z_{j}-g_{n}\right\}\mathbb{I}\left(\mathcal{B}_{j}\right)\mathbb{I}(Z_{j}\geq 0)\right]$
$\geq\exp\left\{-g_{n}\right\}\mathbb{P}_{t_{j}}\left(\mathcal{B}_{j}\cap\left\{Z_{j}\geq 0\right\}\right).$
@@
$\mathbb{P}\left(\mathcal{B}_{j}\cap\left\{Z_{j}\geq 0\right\}\right)=\mathbb{P}(\mathcal{B}_{j})+\mathbb{P}(Z_{j}\geq 0)-\mathbb{P}_{t_{j}}(\mathcal{B}_{j}\cup\left\{Z_{j}\geq 0\right\})\geq\frac{3}{4}+\frac{1}{2}-1=\frac{1}{4}.$
@@
[…]
Thus, $\exp\left\{-g_{n}\right\}n/\log n\to\infty$ for $n\to\infty$, which is a contradiction to (31), so the statement is true.
@@
Thus, $\bigl{(}{n^{1/3}}/{\log^{{1}/{3}}(n)}\bigr{)}\exp\left\{-g_{n}\right\}\to\infty$, for $n\to\infty$, which is a contradiction to (31). ∎